OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for cooking classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 11(1.35)x The average rate of change in the number of students who enrolled for cooking classes from the first to the fifth year is ________students per year. Round your answer to the nearest whole number. (4 points)

OpenStudy (agl202):

Well just solve for f(x) five times (with 1, 2, 3, 4, and 5) and then find the average rate of change, and to do that you just see how much they changed each year, add them all up, and divide by 5.

OpenStudy (anonymous):

so how much would it be?

OpenStudy (agl202):

When rounding u would be getting 6...

OpenStudy (anonymous):

are u sure?

OpenStudy (agl202):

Im sure, gotta go

OpenStudy (anonymous):

ok thanks but can you help only with 1more

OpenStudy (agl202):

If I can...

OpenStudy (anonymous):

The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 2 4 8 32 Which recursive function best represents the values shown in the table? (4 points) f(1) = 1, f(2) = 2, f(n) = f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (agl202):

The second option should be right: if you know that n > 2, f(3), which is three, needs to be f(1) plus f(2). The second option is the only way that represents that correctly

OpenStudy (agl202):

WoW!!!!

OpenStudy (anonymous):

ok thanks you cant help me with 2 more right?

OpenStudy (anonymous):

Simplify. 

OpenStudy (agl202):

Any answer choices...

OpenStudy (anonymous):

mmmm